OpenStudy (anonymous):

Math Analysis: Solve for 0°

OpenStudy (accessdenied):

try substituting u=sin(x) with that, you'd have.. \[ 6u^2 - 7u + 2 = 0 \] you can solve for u by factoring and then back-substitute for equations in the form "sin x = ..."

OpenStudy (anonymous):

it can be (x+3 ) (x+6) =0 or (x+9/2) -9/4 =0

OpenStudy (anonymous):

ok i got u=-2/3 and u=-1/2

OpenStudy (anonymous):

what's next?

OpenStudy (accessdenied):

hmm.. can you show what you did? I get u=2/3 and u=1/2 6u^2 - 4u - 3u + 2 = 0 2u(3u - 2) - (3u - 2) = 0 (3u - 2)(2u - 1) = 0 3u - 2 = 0 2u - 1 = 0 3u = 2 2u = 1 u = 2/3 u = 1/2

OpenStudy (anonymous):

oops nvm i got it thx for ur time! :P

OpenStudy (accessdenied):

No problem. :)